# Hausdorff Dimension of Singular Vectors in Function Fields

### v1

@@ 5. Upper Bound
Fix $\boldsymbol{\theta}\in\widetilde{\mathbf{DI}}_{d}^{*}(\varepsilon)$. We first claim that the sequence of best approximations $\{\mathbf{u}_{n}\}_{n\in\mathbb{N}}$ of $\boldsymbol{\theta}$ contains a $\sigma_{\varepsilon}$-admissible sequence. By definition, there exists some $n_{0}$ such that $\mathbf{u}_{n}\in Q_{\varepsilon}$ for every $n>n_{0}$. Choose $n_{1}=n_{0}+1$ and assume that we have already constructed a $\sigma_{\varepsilon}$-admissible sequence $\mathbf{u}_{n_{1}},\dots,\mathbf{u}_{n_{i}}$. We let $n_{i+1}$ be the smallest $m>n_{i}$ such that $\mathbf{u}_{m}\notin H_{\mathbf{u}_{n_{i}}}$. Since $1,\theta_{1}\dots\theta_{d}$ are linearly independent over $\mathcal{K}$, such $n_{i+1}$ indeed exists. Now we let $\mathbf{u}=\mathbf{u}_{n_{i}}$, $\mathbf{v}=\mathbf{u}_{n_{i+1}-1}$, and $\mathbf{w}=\mathbf{u}_{n_{i+1}}$. Then, $|\mathbf{u}|\leq|\mathbf{v}|<|\mathbf{w}|$, $\mathbf{v}\in H_{\mathbf{u}}$, and $\mathbf{w}\notin H_{\mathbf{u}}$. Moreover, it follows from Lemmas 4.6 and 4.10 that
@@
for any $\mathbf{v}\in D_{k}(\mathbf{u})$. Hence, it follows that $\pi_{\mathbf{u}}(D_{k}(\mathbf{u}))\subseteq B(0,q^{k}\lambda_{1}(\Lambda_{\mathbf{u}}))\cap\Lambda_{\mathbf{u}}^{\prime}$. Thus, by Lemma 4.13, Proposition 2.5, and the fact that $\lambda_{1}(\Lambda_{\mathbf{u}})\leq\lambda_{i}(\Lambda_{\mathbf{u}}^{\prime})$ for every $i=1,\dots,d-1$, we obtain
#$\begin{split}\#D_{k}(\mathbf{u})\leq(q-1)q^{k}\#\left(\Lambda_{\mathbf{u}}^{\prime}\cap B(0,q^{k}\lambda_{1}(\Lambda_{\mathbf{u}}))\right)&=(q-1)q^{k}\prod_{i=1}^{d-1}\Bigg{\lceil}\frac{q^{k+1}\lambda_{1}(\Lambda_{\mathbf{u}})}{\lambda_{i}(\Lambda_{\mathbf{u}}^{\prime})}\Bigg{\rceil}\\ &\leq(q-1)q^{d-1}q^{kd}.\end{split}$ (5.2)
@@
#$\begin{split}\#E_{k}(\mathbf{u},\mathbf{v},\varepsilon)&\leq(q-1)q^{k}\#\left(B(0,q^{\varepsilon k\frac{d-1}{d}}|\mathbf{v}|^{-\frac{1}{d}})\cap\Lambda_{\mathbf{v}}\setminus H_{\mathbf{u}}^{\prime}\right)\\ &=(q-1)q^{k}\prod_{i=1}^{d}\Bigg{\lceil}\frac{q\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-1/d}}{\lambda_{i}(\Lambda_{\mathbf{v}})}\Bigg{\rceil}\\ &\leq(q-1)q^{k}\prod_{i=1}^{d}\left(\frac{q\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-1/d}+\lambda_{d}(\Lambda_{\mathbf{v}})}{\lambda_{i}(\Lambda_{\mathbf{v}})}\right)\\ &\leq(q-1)q^{k}\prod_{i=1}^{d}\frac{(q+q^{2})\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-1/d}}{\lambda_{i}(\Lambda_{\mathbf{v}})}\\ &=(q-1)q^{k}(q+q^{2})^{d}\varepsilon^{d}q^{k(d-1)}.\end{split}$ (5.5)
@@
For any $\frac{d^{2}}{d+1}<s\leq d$, it follows from Propositions 5.3 and 5.4 with $t=\frac{d+1}{d}s$ that
@@ 6.1. Self-similar structure
While a certain orthogonal projection was used in [CC16], we will utilize the orthogonal basis given in Definition 4.9. The basic idea is to directly use the lattice given in [CC16, Lemma 8.7]. For this, we need the following definition.
Given $\mathbf{u}\in Q$, we fix an orthogonal basis $\xi_{1},\dots,\xi_{d}$ of the lattice $\Lambda_{\mathbf{u}}$ with $\|\xi_{i}\|=\lambda_{i}(\mathbf{u})$ for each $i=1,\dots,d$ as in Definition 4.9. Note that by Theorem 2.4, $\det(\Lambda_{\mathbf{u}})=\lambda_{1}(\mathbf{u})\cdots\lambda_{d}(\mathbf{u})=1/|\mathbf{u}|$.
@@
    where $|\mathbf{u}\wedge\mathbf{v}|=|\mathbf{u}||\mathbf{v}|\|\hat{\mathbf{u}}-\hat{\mathbf{v}}\|$. For each $i=1\dots,d-1$, denote $\lambda_{i}(\alpha)=\lambda_{i}(\Lambda_{\alpha})$ and define $\hat{\lambda}_{i}(\alpha)=\hat{\lambda}_{i}(\Lambda_{\alpha})=|\mathbf{u}\wedge\mathbf{v}|^{\frac{1}{d-1}}\lambda_{i}(\Lambda_{\alpha})$.
@@
Now, let $\beta=\pi_{\mathbf{v}}(\mathbf{u})$. For any $\gamma\in\Lambda_{\mathbf{v}}\setminus\widetilde{\mathcal{K}}\beta$, since $\mathrm{Ker}(\pi_{\alpha})=\widetilde{\mathcal{K}}\beta$, it follows from the claim and the observation that $\|\gamma\|\geq\|\pi_{\alpha}(\gamma)\|\geq\lambda_{1}(\alpha)\geq\frac{|\mathbf{u}\wedge\mathbf{v}|}{|\mathbf{v}|}$. Since $\alpha$ is primitive, $\beta$ is also primitive. Hence, if $\gamma\in\Lambda_{\mathbf{v}}\cap\widetilde{\mathcal{K}}\beta\setminus\{0\}$, then $\|\gamma\|\geq\|\beta\|=\frac{|\mathbf{u}\wedge\mathbf{v}|}{|\mathbf{v}|}$. Therefore, we have $\lambda_{1}(\mathbf{v})=\frac{|\mathbf{u}\wedge\mathbf{v}|}{|\mathbf{v}|}$. ∎
@@
We need to verify that for every $\sigma_{\varepsilon,N}$ admissible sequence $\{\mathbf{u}_{k}\}_{k\in\mathbb{N}}$,
@@ 6.3. Lower bound calculations
Let $(Q_{\varepsilon,N},\sigma_{\varepsilon,N},B)$ be the triple given in Definition 6.4, which is a strictly nested self-similar structure covering a subset of $\mathbf{DI}_{d}(\varepsilon)$ by Proposition 6.7. In order to use Theorem 3.3, we first verify the condition (1) of Theorem 3.3. Indeed, since $\hat{\lambda}_{d}(\mathbf{u})\geq 1$, it follows from Lemma 6.10 that (1) of Theorem 3.3 holds with $c=0$ and $\rho=q\left(\frac{\varepsilon}{q^{N+1}}\right)^{\frac{2d}{d-1}}$.
Fix small $\delta>0$ and let
@@
Now, we claim that $\left(\frac{\varepsilon}{q}\right)^{\frac{d^{2}}{d-1}f(\varepsilon)}\geq q^{-\frac{d^{2}}{d-1}}$. Indeed, it is enough to show that $f(\varepsilon)\log_{q}\frac{q}{\varepsilon}\leq 1$ and since $\log(1+y)\leq y$ and $\log y\leq y$ for any $y>0$, we have
@@
Let $\mathbf{u}_{-1}\in Q$ and let $\mathbf{u}_{0}\in F_{N_{0}}(\mathbf{u}_{-1},\varepsilon_{0})$. Define $Q_{0}=\{\mathbf{u}_{0}\}$ and $Q_{i+1}=\bigcup_{\mathbf{u}\in Q_{i}}F_{N_{i}}(\mathbf{u},\varepsilon_{i})$ for each $i\geq 0$. For each $\mathbf{u}\in Q_{i}$, define $\sigma^{\prime}(\mathbf{u})=F_{N_{i}}(\mathbf{u},\varepsilon_{i})$ and $B(\mathbf{u})=B\left(\hat{\mathbf{u}},\frac{\lambda_{1}(\mathbf{u})}{q|\mathbf{u}|}\right)$. Finally, let $Q^{\prime}=\bigcup_{i\geq 0}Q_{i}$ be a disjoint union. By the same arguments of the proof of Proposition 6.7, the triple $(Q^{\prime},\sigma^{\prime},B)$ is a strictly nested self-similar structure covering a subset of $\mathbf{Sing}_{d}$.
By Lemma 6.10, for $\mathbf{u}\in Q_{i}$ and any distinct $\mathbf{v},\mathbf{w}\in F_{N_{i}}(\mathbf{u},\varepsilon_{i})$ we have
$d\left(B(\mathbf{v}),B(\mathbf{w})\right)\geq\left(\frac{\varepsilon_{i}}{q^{N_{i}+1}\hat{\lambda}_{d}(\mathbf{u})}\right)^{\frac{2d}{d-1}}\frac{\lambda_{1}(\mathbf{u})}{|\mathbf{u}|}\geq q\left(\frac{\varepsilon_{i}}{q^{N_{i}+1}}\right)^{\frac{2d}{d-1}}\frac{\lambda_{1}(\mathbf{u})}{q|\mathbf{u}|}.$
@@
$\rho(\mathbf{u})=q\left(\frac{\varepsilon_{i}}{q^{N_{i}+1}}\right)^{\frac{2d}{d-1}}\quad\text{for each }\mathbf{u}\in Q_{i}.$
@@
$\begin{split}\sum_{\mathbf{v}\in\sigma^{\prime}(\mathbf{u})}\frac{(\rho(\mathbf{v})\operatorname{diam}B(\mathbf{v}))^{s}}{(\rho(\mathbf{u})\operatorname{diam}B(\mathbf{u}))^{s}}&=\left(\frac{\varepsilon_{i+1}}{\varepsilon_{i}q^{N_{i+1}-N_{i}}}\right)^{\frac{2d}{d-1}s}\sum_{\mathbf{v}\in F_{N_{i}}(\mathbf{u},\varepsilon_{i})}\left(\frac{\lambda_{1}(\mathbf{v})}{|\mathbf{v}|}\right)^{s}\left(\frac{|\mathbf{u}|}{\lambda_{1}(\mathbf{u})}\right)^{s}\\ &\geq\left(\frac{\varepsilon_{i+1}}{\varepsilon_{i}q}\right)^{\frac{2d}{d-1}s}\left(\frac{q-1}{q}-\varepsilon_{i}^{d-1}\right)\frac{\varepsilon_{i}^{d}(q-1)^{2}}{q^{d+s}}\left(\frac{d}{d-1}N_{i}-2\right).\end{split}$
@@
As in the proofs of Propositions 5.3 and 5.4, we will divide $F_{N}(\mathbf{u},\varepsilon)$ by the following subsets
@@
In order to prove Proposition 6.15, we follow the strategy of [CC16, Section 8]. Let $\mathbf{u}\in Q$ and $\varepsilon>0$ be given. For any nonzero $n\in\mathcal{R}$, we will first give an lower bound of #$\#(\Lambda_{\mathbf{u}}(\varepsilon)\cap C_{n}^{\prime}(\mathbf{u}))$. By definition, we have
@@
#$\begin{split}\#(\Lambda_{\mathbf{u}}(\varepsilon)\cap C_{n}^{\prime}(\mathbf{u}))&\geq\#X_{n}\times\#\Bigg{\{}\mathbf{k}\in\mathcal{R}^{d-1}:\forall i,\ |k_{i}|\leq\frac{\lambda_{d}(\mathbf{u})}{\lambda_{i}(\mathbf{u})}\Bigg{\}}\\ &=\#X_{n}\times\frac{\lambda_{d}^{d-1}(\mathbf{u})}{\lambda_{1}(\mathbf{u})\cdots\lambda_{d-1}(\mathbf{u})}\end{split}$ (6.10)

### v2

@@ 5. Upper Bound
Remark 5.2.
We will consider the triple $(Q_{\varepsilon},\sigma_{\varepsilon},B)$ as a self-similar structure on $\widetilde{\mathcal{K}}^{d}$. Recall Definition 3.1 and note that $(\mathbf{u},\mathbf{w})\in\sigma_{\varepsilon}$ if and only if $\mathbf{w}\in E(\mathbf{u},\mathbf{v},\varepsilon)$ for some $\mathbf{v}\in D(\mathbf{u})$. Observe that $\sigma_{\varepsilon}(\mathbf{u})\subseteq Q_{\varepsilon}$ for every $\mathbf{u}\in Q$ since for any $\mathbf{w}\in\sigma_{\varepsilon}(\mathbf{u})$ we have
$\hat{\lambda}_{1}(\mathbf{w})\leq|\mathbf{w}|^{\frac{1}{d}}\|\pi_{\mathbf{w}}(\mathbf{v})\|=|\mathbf{w}|^{\frac{1}{d}}|\mathbf{v}|\cdot\|\hat{\mathbf{v}}-\hat{\mathbf{w}}\|<\varepsilon.$
Thus, the triple $(Q_{\varepsilon},\sigma_{\varepsilon},B)$ is indeed a self-similar structure.
@@
Fix $\boldsymbol{\theta}\in\widetilde{\mathbf{DI}}_{d}^{*}(\varepsilon)$. We now claim that the sequence of best approximations $\{\mathbf{u}_{n}\}_{n\in\mathbb{N}}$ of $\boldsymbol{\theta}$ contains a $\sigma_{\varepsilon}$-admissible sequence. By definition, there exists some $n_{0}$ such that $\mathbf{u}_{n}\in Q_{\varepsilon}$ for every $n>n_{0}$. Choose $n_{1}=n_{0}+1$ and assume that we have already constructed a $\sigma_{\varepsilon}$-admissible sequence $\mathbf{u}_{n_{1}},\dots,\mathbf{u}_{n_{i}}$. We let $n_{i+1}$ be the smallest $m>n_{i}$ such that $\mathbf{u}_{m}\notin H_{\mathbf{u}_{n_{i}}}$. Since $1,\theta_{1}\dots\theta_{d}$ are linearly independent over $\mathcal{K}$, such $n_{i+1}$ indeed exists. Now we let $\mathbf{u}=\mathbf{u}_{n_{i}}$, $\mathbf{v}=\mathbf{u}_{n_{i+1}-1}$, and $\mathbf{w}=\mathbf{u}_{n_{i+1}}$. Then, $|\mathbf{u}|\leq|\mathbf{v}|<|\mathbf{w}|$, $\mathbf{v}\in H_{\mathbf{u}}$, and $\mathbf{w}\notin H_{\mathbf{u}}$. Moreover, it follows from Lemmas 4.6 and 4.10 that
@@
for any $\mathbf{v}\in D_{k}(\mathbf{u})$. Hence, it follows that $\pi_{\mathbf{u}}(D_{k}(\mathbf{u}))\subseteq B(0,q^{k}\lambda_{1}(\Lambda_{\mathbf{u}}))\cap\Lambda_{\mathbf{u}}^{\prime}$.
Let $\xi_{1},\dots,\xi_{d}$ be an orthogonal basis of the lattice $\Lambda_{\mathbf{u}}$ with $\|\xi_{i}\|=\lambda_{i}(\Lambda_{\mathbf{u}})$ for each $i=1,\dots,d$ as in the paragraph after Definition 4.8. Since $\Lambda_{\mathbf{u}}^{\prime}=\mathrm{span}_{\mathcal{R}}\{\xi_{1},\dots,\xi_{d-1}\}$, it follows that
#$\begin{split}\#\left(B(0,q^{k}\lambda_{1}(\Lambda_{\mathbf{u}}))\cap\Lambda_{\mathbf{u}}^{\prime}\right)&=\#\left\{\sum_{i=1}^{d-1}a_{i}\xi_{i}:a_{i}\in\mathcal{R},|a_{i}|\leq q^{k}\frac{\lambda_{1}(\Lambda_{\mathbf{u}})}{\lambda_{i}(\Lambda_{\mathbf{u}})}\right\}\\ &\leq\prod_{i=1}^{d-1}\#\{a_{i}\in\mathcal{R}:|a_{i}|\leq q^{k}\}=q^{(k+1)(d-1)}.\end{split}$
Thus, it follows from Lemma 4.13 that
#$\#D_{k}(\mathbf{u})\leq(q-1)q^{k}\#\left(B(0,q^{k}\lambda_{1}(\Lambda_{\mathbf{u}})\cap\Lambda_{\mathbf{u}}^{\prime})\right)\leq(q-1)q^{d-1}q^{kd}.$ (5.2)
@@
#$\begin{split}\#E_{k}(\mathbf{u},\mathbf{v},\varepsilon)&\leq(q-1)q^{k}\#\left(B(0,\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-\frac{1}{d}})\cap\Lambda_{\mathbf{v}}\setminus H_{\mathbf{u}}^{\prime}\right)\\ &\leq(q-1)q^{k}\prod_{i=1}^{d}\Bigg{\lceil}\frac{q\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-1/d}}{\lambda_{i}(\Lambda_{\mathbf{v}})}\Bigg{\rceil}\\ &\leq(q-1)q^{k}\prod_{i=1}^{d}\left(\frac{q\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-1/d}+\lambda_{d}(\Lambda_{\mathbf{v}})}{\lambda_{i}(\Lambda_{\mathbf{v}})}\right)\\ &\leq(q-1)q^{k}\prod_{i=1}^{d}\frac{(q+q^{2})\varepsilon q^{k\frac{d-1}{d}}|\mathbf{v}|^{-1/d}}{\lambda_{i}(\Lambda_{\mathbf{v}})}\\ &=(q-1)q^{k}(q+q^{2})^{d}\varepsilon^{d}q^{k(d-1)}.\end{split}$ (5.5)
@@
For any $\frac{d^{2}}{d+1}<s\leq d$, it follows from Propositions 5.4 and 5.5 with $t=\frac{d+1}{d}s$ that
@@ 6.1. Self-similar structure
While a certain orthogonal projection was used in [CC16], we will utilize the orthogonal basis given in the paragraph after Definition 4.8. The basic idea is to directly use the lattice given in [CC16, Lemma 8.7]. For this, we need the following definition.
Given $\mathbf{u}\in Q$, we fix an orthogonal basis $\xi_{1},\dots,\xi_{d}$ of the lattice $\Lambda_{\mathbf{u}}$ with $\|\xi_{i}\|=\lambda_{i}(\mathbf{u})$ for each $i=1,\dots,d$ as in the paragraph after Definition 4.8. Note that by Theorem 2.4, $\det(\Lambda_{\mathbf{u}})=\lambda_{1}(\mathbf{u})\cdots\lambda_{d}(\mathbf{u})=1/|\mathbf{u}|$.
@@
    where $|\mathbf{u}\wedge\mathbf{v}|=|\mathbf{u}||\mathbf{v}|\|\hat{\mathbf{u}}-\hat{\mathbf{v}}\|$. It is worth noting that the quantity $|\mathbf{u}\wedge\mathbf{v}|$ is the supremum norm of the natural projection of $\mathbf{u}\wedge\mathbf{v}\in\bigwedge^{2}\widetilde{\mathcal{K}}^{d+1}$ onto the subspace spanned by $\mathbf{e}_{1}\wedge\mathbf{e}_{d+1},\dots,\mathbf{e}_{d}\wedge\mathbf{e}_{d+1}$. For each $i=1\dots,d-1$, denote $\lambda_{i}(\alpha)=\lambda_{i}(\Lambda_{\alpha})$ and define $\hat{\lambda}_{i}(\alpha)=\hat{\lambda}_{i}(\Lambda_{\alpha})=|\mathbf{u}\wedge\mathbf{v}|^{\frac{1}{d-1}}\lambda_{i}(\Lambda_{\alpha})$.
@@
Now, let $\beta=\pi_{\mathbf{v}}(\mathbf{u})$. For any $\gamma\in\Lambda_{\mathbf{v}}\setminus\widetilde{\mathcal{K}}\beta$, since $\mathrm{Ker}(\pi_{\alpha})=\widetilde{\mathcal{K}}\beta$, it follows from the claim and the observation that $\|\gamma\|\geq\|\pi_{\alpha}(\gamma)\|\geq\lambda_{1}(\alpha)\geq\frac{|\mathbf{u}\wedge\mathbf{v}|}{|\mathbf{v}|}$. Finally, we claim that $\beta$ is primitive. Assuming this claim, if $\gamma\in\Lambda_{\mathbf{v}}\cap\widetilde{\mathcal{K}}\beta\setminus\{0\}$, then $\|\gamma\|\geq\|\beta\|=\frac{|\mathbf{u}\wedge\mathbf{v}|}{|\mathbf{v}|}$. Therefore, $\lambda_{1}(\mathbf{v})=\frac{|\mathbf{u}\wedge\mathbf{v}|}{|\mathbf{v}|}$.
In order to prove the last claim, suppose that $\beta$ is not primitive and write $\beta=f\gamma$ with $\gamma\in\Lambda_{\mathbf{v}}$ and $f\in\mathcal{R}$ with $\mathrm{deg}f\geq 1$. Choose $\mathbf{w}\in Q$ such that $\gamma=\pi_{\mathbf{v}}(\mathbf{w})$. By linearity of $\pi_{\mathbf{v}}$, $\pi_{\mathbf{v}}(f\mathbf{w}-\mathbf{u})=0$, hence $f\mathbf{w}-\mathbf{u}=f^{\prime}\mathbf{v}$ for some $f^{\prime}\in\mathcal{R}$. This implies that $f^{\prime}\alpha=f\pi_{\mathbf{u}}(\mathbf{w})$. Since $\alpha$ is primitive, we have $f|f^{\prime}$. This contradicts the fact that $f\mathbf{w}-f^{\prime}\mathbf{v}=\mathbf{u}\in Q$. ∎
@@
Note that by the definition of $Q_{\varepsilon,N}$, if $\mathbf{u}\in Q_{\varepsilon,N}$, then, $\sigma_{\varepsilon,N}(\mathbf{u})\subseteq Q_{\varepsilon,N}$. We need to verify that for every $\sigma_{\varepsilon,N}$ admissible sequence $\{\mathbf{u}_{k}\}_{k\in\mathbb{N}}$,
@@ 6.3. Lower bound calculations
Let $(Q_{\varepsilon,N},\sigma_{\varepsilon,N},B)$ be the triple given in Definition 6.4, which is a strictly nested self-similar structure covering a subset of $\mathbf{DI}_{d}(\varepsilon)$ by Proposition 6.7. In order to use Theorem 3.3, we first verify the condition (1) of Theorem 3.3. Indeed, recall that $\hat{\lambda}_{d}(\mathbf{u})\leq\frac{q^{d-1}}{\varepsilon^{d-1}}$ for any $\mathbf{u}\in Q_{\varepsilon,N}$ by Remark 6.12, hence it follows from Lemma 6.10 that (1) of Theorem 3.3 holds with $c=0$ and $\rho=q\left(\frac{\varepsilon^{d}}{q^{N+d}}\right)^{\frac{2d}{d-1}}$. Fix small $\delta>0$ and let
@@
We note that since $f(\varepsilon)\leq 1$, $s=\frac{d^{2}}{d+1}+\frac{d}{d+1}f(\varepsilon)\leq d$, hence, the third line of the above equation follows. Now, we claim that $\left(\frac{\varepsilon}{q}\right)^{\frac{d^{2}}{d-1}f(\varepsilon)}\geq q^{-\frac{d^{2}}{d-1}}$. Indeed, it is enough to show that $f(\varepsilon)\log_{q}\frac{q}{\varepsilon}\leq 1$ and since $\log(1+y)\leq y$ and $\log y\leq y$ for any $y>0$, we have
@@
Let $\mathbf{u}_{-1}\in Q$ and let $\mathbf{u}_{0}\in F_{N_{0}}(\mathbf{u}_{-1},\varepsilon_{0})$. Define $Q_{0}=\{\mathbf{u}_{0}\}$ and $Q_{i+1}=\bigcup_{\mathbf{u}\in Q_{i}}F_{N_{i}}(\mathbf{u},\varepsilon_{i})$ for each $i\geq 0$. For each $\mathbf{u}\in Q_{i}$, define $\sigma^{\prime}(\mathbf{u})=F_{N_{i}}(\mathbf{u},\varepsilon_{i})$ and $B(\mathbf{u})=B\left(\hat{\mathbf{u}},\frac{\lambda_{1}(\mathbf{u})}{q|\mathbf{u}|}\right)$. Finally, let $Q^{\prime}=\bigcup_{i\geq 0}Q_{i}$. By the same arguments of the proof of Proposition 6.7, the triple $(Q^{\prime},\sigma^{\prime},B)$ is a strictly nested self-similar structure covering a subset of $\mathbf{Sing}_{d}$.
It follows from Remark 6.12 that $\hat{\lambda}_{d}(\mathbf{u})\leq\frac{q^{d-1}}{\varepsilon_{i-1}^{d-1}}<\frac{q^{d-1}}{\varepsilon_{i}^{d-1}}$ for any $\mathbf{u}\in Q_{i}$. By Lemma 6.10, for $\mathbf{u}\in Q_{i}$ and any distinct $\mathbf{v},\mathbf{w}\in F_{N_{i}}(\mathbf{u},\varepsilon_{i})$ we have
$d\left(B(\mathbf{v}),B(\mathbf{w})\right)\geq\left(\frac{\varepsilon_{i}}{q^{N_{i}+1}\hat{\lambda}_{d}(\mathbf{u})}\right)^{\frac{2d}{d-1}}\frac{\lambda_{1}(\mathbf{u})}{|\mathbf{u}|}>q\left(\frac{\varepsilon_{i}^{d}}{q^{N_{i}+d}}\right)^{\frac{2d}{d-1}}\frac{\lambda_{1}(\mathbf{u})}{q|\mathbf{u}|}.$
@@
$\rho(\mathbf{u})=q\left(\frac{\varepsilon_{i}^{d}}{q^{N_{i}+d}}\right)^{\frac{2d}{d-1}}\quad\text{for each }\mathbf{u}\in Q_{i}.$
@@
$\begin{split}\sum_{\mathbf{v}\in\sigma^{\prime}(\mathbf{u})}\frac{(\rho(\mathbf{v})\operatorname{diam}B(\mathbf{v}))^{s}}{(\rho(\mathbf{u})\operatorname{diam}B(\mathbf{u}))^{s}}&=\left(\frac{\varepsilon_{i+1}^{d}}{\varepsilon_{i}^{d}q^{N_{i+1}-N_{i}}}\right)^{\frac{2d}{d-1}s}\sum_{\mathbf{v}\in F_{N_{i}}(\mathbf{u},\varepsilon_{i})}\left(\frac{\lambda_{1}(\mathbf{v})}{|\mathbf{v}|}\right)^{s}\left(\frac{|\mathbf{u}|}{\lambda_{1}(\mathbf{u})}\right)^{s}\\ &\geq\left(\frac{\varepsilon_{i+1}^{d}}{\varepsilon_{i}^{d}q}\right)^{\frac{2d}{d-1}s}\left(\frac{q-1}{q}-\varepsilon_{i}^{d-1}\right)\frac{\varepsilon_{i}^{d}(q-1)^{2}}{q^{d+s}}\left(\frac{d}{d-1}N_{i}-2\right).\end{split}$
@@
As in the proofs of Propositions 5.4 and 5.5, we will divide $F_{N}(\mathbf{u},\varepsilon)$ into the following subsets
@@
In order to prove Proposition 6.15, we follow the strategy of [CC16, Section 8]. Let $\mathbf{u}\in Q$ and $\varepsilon>0$ be given. For any nonzero $n\in\mathcal{R}$, we will first give a lower bound of #$\#(\Lambda_{\mathbf{u}}(\varepsilon)\cap C_{n}^{\prime}(\mathbf{u}))$. By definition, we have
@@
#$\begin{split}\#(\Lambda_{\mathbf{u}}(\varepsilon)\cap C_{n}^{\prime}(\mathbf{u}))&\geq\#X_{n}\times\#\Bigg{\{}\mathbf{k}\in\mathcal{R}^{d-1}:\forall i,\ |k_{i}|\leq\frac{\lambda_{d}(\mathbf{u})}{\lambda_{i}(\mathbf{u})}\Bigg{\}}\\ &=\#X_{n}\times\frac{\lambda_{d}^{d-1}(\mathbf{u})}{\lambda_{1}(\mathbf{u})\cdots\lambda_{d-1}(\mathbf{u})}.\end{split}$ (6.10)